OpenStudy (anonymous):

Suppose you roll two number cubes and pick a letter of the alphabet at random. Find the probability you roll 2 even numbers and pick one of the vowels a, e, i, o, or u. A. 7/104 B. 5.104 C. 1/6 D. 1/104

jimthompson5910 (jim_thompson5910):

Hint: assuming the events are independent, you would multiply out the probabilities

jimthompson5910 (jim_thompson5910):

P(even number, even number, vowel) = P(even number)*P(even number)*P(vowel)

OpenStudy (anonymous):

its 1/104

jimthompson5910 (jim_thompson5910):

P(even number) = ???

jimthompson5910 (jim_thompson5910):

There are 3 even numbers (2,4,6) Out of 6 total (1,2,3,4,5,6) So... P(even number) = ???

OpenStudy (anonymous):

oh its 1/12

OpenStudy (anonymous):

wait that's not right

OpenStudy (anonymous):

even number is 12

jimthompson5910 (jim_thompson5910):

if you roll the die, what's the probability of rolling an even number?

jimthompson5910 (jim_thompson5910):

keep in mind that 3 out of the 6 numbers are even

jimthompson5910 (jim_thompson5910):

it's not 1/12

OpenStudy (anonymous):

so then its 1/6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you all :)

jimthompson5910 (jim_thompson5910):

P(even number) = 3/6 P(even number) = 1/2

jimthompson5910 (jim_thompson5910):

P(vowel) = 5/26

jimthompson5910 (jim_thompson5910):

P(even number, even number, vowel) = P(even number)*P(even number)*P(vowel) P(even number, even number, vowel) = (1/2)*(1/2)*(5/26) I'll let you multiply it out The answer is NOT 1/6